OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 8x + 2y = 4 –4x + 3y = –10 help??

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OpenStudy!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

Have you used elmination before?

OpenStudy (kidrah69):

@Trevorhowell16

OpenStudy (anonymous):

no im behin in school ialready got that tho

OpenStudy (kidrah69):

Ok so we need opposite signs so that the X's will cancel out :) \[2(–4x + 3y = –10)\] multiply all by -2 let me know what your new equation is :)

OpenStudy (anonymous):

im done with it haha

OpenStudy (kidrah69):

oh xD sorry so you dont need help? :)

OpenStudy (anonymous):

Elena bought 320 tropical fish for a museum display. She bought 7 times as many triggerfish as parrotfish. How many of each type of fish did she buy? Which is a system of equations to model the problem if x represents the number of parrotfish Elena bought and y represents the number of triggerfish she bought? (Points : 1) x + 7y = 320 y = 7x x + y = 7 y = 320x x + y = 320 y = 7x x – y = 320 y = 7x

OpenStudy (anonymous):

>

OpenStudy (nincompoop):

assign a variable for triggerfish and for parrotfish

OpenStudy (anonymous):

dont now how

OpenStudy (anonymous):

no

OpenStudy (nincompoop):

a variable is any letter or symbol assigned for a particular object so if you're the object, I can use T

OpenStudy (nincompoop):

if the object we are looking at is an apple, I can assign a variable a